# FastRAG: Retrieval Augmented Generation for Semi-structured Data

### v1

@@ Abstract
Efficiently processing and interpreting network data is critical for the operation of increasingly complex networks. Recent advances in Large Language Models (LLM) and Retrieval-Augmented Generation (RAG) techniques have improved data processing in network management. However, existing RAG methods like VectorRAG and GraphRAG struggle with the complexity and implicit nature of semi-structured technical data, leading to inefficiencies in time, cost, and retrieval.
This paper introduces FastRAG, a novel RAG approach designed for semi-structured data. FastRAG employs schema learning and script learning to extract and structure data without needing to submit entire data sources to an LLM. It integrates text search with knowledge graph (KG) querying to improve accuracy in retrieving context-rich information. Evaluation results demonstrate that FastRAG provides accurate question answering, while improving up to $90\%$ in time and $85\%$ in cost compared to GraphRAG.
@@ I Introduction
Efficiently processing and understanding semi-structured data is crucial to improve network management tasks. As networks grow, the diversity of platforms and the sheer volume of data introduce significant challenges. Traditional tools for processing network data [1, 2, 3] offer some level of utility but fall short in comprehensively extracting and utilizing the full range of information embedded in semi-structured data formats such as logs and configurations. Moreover, the task of correlating these data from different vendors is further complicated by disaggregated network service implementation. Although natural language processing (NLP) and Machine Learning (ML) algorithms have been recently developed to process these data [4, 5], these solutions often lack the flexibility needed to handle diverse data types and formats across varying network environments.
@@
The emergence of Large Language Models (LLMs) has introduced new possibilities for processing and understanding natural language, showcasing their potential in improving network management tools. However, using pre-trained models for generating domain-specific and coherent responses remains challenging. Techniques like Retrieval-Augmented Generation (RAG) [6] have been introduced to enhance LLMs by integrating retrieval methods, with VectorRAG providing context retrieval from textual documents based on semantic similarity [7]. While effective in many scenarios, these traditional RAG systems have limitations, such as loss of critical contextual information due to uniform chunking of documents [8]. GraphRAG [9], a more advanced approach, addresses some of these limitations by leveraging knowledge graphs (KG) to organize information extracted from source documents. Despite its advantages, GraphRAG struggles with queries for accurate information or lacking explicit entities. To overcome these challenges, HybridRAG [8] combines the strengths of both VectorRAG and GraphRAG, offering improved accuracy in retrieving relevant information for LLMs.
@@
Despite these advancements, existing RAG techniques have limitations when applied to network data. The reliance on embedding vectors for context retrieval does not perform well with semi-structured technical data, where implicit information may be hidden behind specific keywords. Moreover, current RAG systems rely on LLMs to extract structured information from source documents by processing them chunk-by-chunk, resulting in increased processing time and cost.
To address these issues, this paper explores a novel RAG approach designed for a cost-effective processing of the large volumes of semi-structured data generated by modern networks. The proposed system (FastRAG) processes source data without requiring the submission of all chunks through an LLM. The contributions of this paper include: the development of prompting methods that use schema learning and script learning to process and structure source data, an algorithm to select sample chunks for these learning prompts, and a retrieval approach that integrates text search and graph querying via Graph Query Language (GQL) for more accurate information retrieval in question answering tasks.
@@
The rest of this paper is organized as follows: Section II reviews related work and discusses the motivation behind developing FastRAG. Section III presents the detailed design of the FastRAG system, including its key components and techniques. The evaluation of FastRAG is discussed in Section IV. Finally, Section V outlines the system’s limitations and concludes the paper.
@@ II Related Work and Motivation
The integration of LLMs into network management has opened new possibilities, particularly in automating network configuration tasks. Recent studies have explored LLMs’ potential to generate network configurations [10, 11, 12], detect anomalies [13], and convert textural descriptions into formal specifications [14, 15]. These contributions represent the most frequent applications of LLMs to network management, aiming to reduce manual labor and risks of errors. The work proposed in this paper is complementary to these approaches as they may rely on RAG systems to retrieve network-specific information.
@@
The application of RAG within network management has been less explored. RAGLog [16] introduces a novel method for detecting log anomalies using a combination of vector databases and LLMs. This approach allows for the anomalies based on raw log data, adapting to various log sources without extensive preprocessing. While promising for its adaptability, RAGLog encounters challenges related to resource consumption and latency, which could limit its scalability in larger network data. Telco-RAG [17] targets the processing of 3GPP documents. A pipeline with query enhancement demonstrates improvements in the accuracy of LLMs when handling complex questions related to the telecommunication domain.
@@
More generally, VectorRAG, the predominant approach in RAG systems, computes vector embeddings for fixed-size chunks and uses semantic similarity to retrieve information relevant to the input query. However, this method struggles with the nature of network data where domain-specific keywords have different meaning in their context. GraphRAG [9] is a more advanced approach that integrates KGs with text chunking and embedding. The LLM is used to extract entities and relationships and summarize communities detected in the KG, providing a more structured retrieval (also uses semantic similarity). Although GraphRAG is efficient in summarizing and reporting tasks (see Figure 1(a)), it falls short in delivering useful answers, particularly when the query includes specific values, such as names or types, that need to be matched exactly in the retrieval as reported in Figure 1(b). HybridRAG [8] combines GraphRAG and VectorRAG. Although this hybrid approach enhances the accuracy and contextual relevance of responses, it inherits the limitations of its base approaches. Another approach uses entity-relationship extraction to create a KG capturing the important information of the source data. The KG is then searched like a database by translating the textural query into a database query [18]. While this allows for accurate retrieval, it can be impractical given the imperfect parsing obtained with LLMs; as shown in Figure 2 where we tested the accuracy of LLM-based configuration parsing against Batfish [2].
@@
To address these limitations, FastRAG extracts entities and their properties from source data while mapping each entity to specific lines of the original text. This method enables accurate retrieval using a KG while allowing for text search to handle vague queries and compensate imperfect entity extraction. Text search, unlike semantic similarity, matches text based on exact wording or structure [19]. Both KG and text searches are performed using GQL queries on a single KG implemented in a graph database. Unlike other methods, FastRAG avoids costly processing of all source data through LLMs by generating JSON schemas and Python code for data structuring and parsing. While demonstrated on logs and configuration data, this approach can be applied to other semi-structured data, such as playbooks and alarms. To the best of our knowledge, this is the first RAG system to rely solely on code generation for data processing.
(a) Good answer for information summary
(b) Wrong answer for accurate question
Figure 1: Example questions and answer with GraphRAG
Figure 2: Accuracy of an LLM in parsing configuration files
@@ III FastRAG Design
We designed a hybrid approach for both information extraction and retrieval. Instead of using indiscriminate chunking or user-defined schema, our method relies on converting data into a simple automatically-generated JSON structure, while maintaining a link to the original data in a KG [20]. The RAG further relies on LLMs to interact with the KG, utilizing GQL query generation to fetch and interpret information.
Given the extensive and continually updating nature of network data, minimizing processing time and cost is essential. To achieve this, we leverage prompt engineering findings which indicate that LLMs are more effective at generating code rather than directly extracting information into structured formats. Consequently, we introduce schema learning and script learning techniques, where we sample the source data and iteratively task the LLM to generate JSON schemas to structure the information and Python code to extract entities.
An overview of the FastRAG architecture is illustrated in Figure 3 and its components are discussed below.
Figure 3: FastRAG architecture
@@ III-A Chunk sampling
Even with expansive context windows of 100k to 200k tokens, LLMs may struggle in processing large data within a single prompt due to context retention limitations [21]. As a result, it is common practice to divide the data into small chunks, typically of uniform size with an overlap between consecutive chunks, and submit these chunks to the LLM through multiple calls. However, submitting all the chunks can be inefficient. To address this, we exploit the repetitive formatting present in semi-structured data, selecting a representative set of chunks—or sample chunks—that effectively covers the syntax of the source data [13].
The chunk sampling process involves two key procedures: keyword extraction and chunk selection.
@@ III-A1 Keyword extraction
We draw inspiration from NLP techniques to identify the most meaningful terms in the source data [4]. The process begins with the preprocessing of the text, where punctuation is removed and the content is tokenized into individual words. Given that English is the predominant language used in most semi-structured network data, it is the language adopted for this process. After preprocessing, the text is split into lines, and a matrix representation is created based on word frequency, where each row corresponds to a line and each column to a unique term from the entire text. The matrix, which reflects the frequency of each term in each line, serves as the input for the K-means clustering algorithm. This algorithm groups the lines into $n_{c}$ number of clusters based on the similarity of their term frequency patterns. Within each cluster, the $n_{t}$ closest terms to the centroids are selected as keywords. The output is a set of keywords that encapsulate the primary terms present in the text corpus.
@@ III-A2 Samples selection
We designed an algorithm that selects the smallest set of chunks that collectively contain the complete set of extracted keywords. However, merely adding chunks until all keywords are covered does not guarantee the selection of the (near) minimum number of sample chunks. To address this, the algorithm selects chunks based on their entropy that provides an estimate of the diversity introduced by each chunk.
The procedure (see Algorithm 1) begins with the preprocessing of the chunks (as in the previous step) where only the extracted keywords are retained. The preprocessed chunks are then used to compute Term Frequency-Inverse Document Frequency (TF-IDF) vectors, representing the importance of terms within each chunk. Next, the Shannon entropy for each chunk is calculated to gauge the informational diversity it contributes. The algorithm iteratively selects a subset of chunks to maximize term coverage, focusing on those that introduce new terms and the most diversity, until the desired coverage threshold—fixed at 1 in all cases—is reached.
@@
The number of sample chunks selected is directly influenced by the number of extracted keywords, which can be adjusted by varying the parameters $n_{c}$ (number of clusters) and $n_{t}$ (number of terms). This process is discussed in greater detail in Section IV.
@@ III-B Schema learning
Building on the selected sample chunks, we develop a prompt strategy that guides the LLM in identifying entity types and their properties, focusing on schema extraction rather than specific entities. The process (Figure 5) begins with the first sample chunk, where the LLM is prompted to identify and structure the entity types and properties into a JSON schema. In subsequent prompts, the previous JSON schema is iteratively refined by submitting new chunks to the LLM. To avoid complexity, the LLM is instructed to limit high levels of nesting and refrain from creating objects with an excessive number of properties.
After each prompt, the strategy includes a verification step to ensure that the JSON schema generated by the LLM is well-formed. If the schema contains errors, the LLM is asked to correct it based on the provided error message. Each prompt call can be retried up to $N=4$ times.
From the finalized schema, two types of objects are extracted, as illustrated in Figure 4. The first is the Step 1 schema, which is a JSON schema that includes only the level-1 entity types—referred to as sections. Each entity type includes a description [^1] Not shown in the Figure and a string property that contains the source data lines defining the objects belonging to that section. The second, Step 2 schema, is a map where each section is associated with an array where objects correspond to the full schema of that section. This division into Step 1 and Step 2 schemas allows for a more structured approach to data extraction, as discussed below.
@@ III-C Script learning
Similar to the approach used in schema learning, the script learning prompt strategy (Figure 5) begins by instructing the LLM to generate parsing functions based on the first sample chunk. The LLM is then prompted iteratively to refine the previous function code by submitting new sample chunks. To ensure both the syntactic and functional correctness of the generated code, the prompt includes a verification step where the code is executed with the sample chunk used to generate it. If the code contains errors, the LLM is asked to correct them based on the feedback message. Similarly, each prompt call can be retired up to N times.
In the Step 1, the prompt instructs the LLM to map each line in the source data to its corresponding section according to the schema. Once the function is generated, it is used to parse the entire source data and divide it into sections. After that, each section is further split into fixed-size chunks, and the sampling process is repeated for each section. At this stage, each section has its own schema (contained in the Step 2 schema object) and a set of sample chunks, enabling it to be processed independently, following the same method as in Step 1.
In the Step 2, script learning is utilized to generate a specific parsing function for each section. Once these functions are generated, the data within each section is processed using the corresponding parsing function. Processing sections independently enables the LLM to concentrate more effectively on each entity type and also allows for a more targeted refinement of the sample chunks by re-executing the sampling process for each section.
@@ III-D Information Retrieval
The extraction process results in a JSON object where each entity is represented, and within this object, an input_data property contains the lines from the source data that define the entity. This JSON structure serves as the foundation for the KG creation and the information retrieval.
@@ III-D1 KG Creation
Each entity identified in the JSON object is inserted as a node within the KG, with the entity’s type serving as the node’s label. Simple-type properties of the entity are directly assigned as properties of the node, while properties that are themselves objects are inserted as child nodes, connected to their parent node. Additionally, for each line within the input_data property of each entity, a corresponding node is created and linked to the parent entity node. These input data nodes are used for text searches leveraging NLP methods [22].
@@ III-D2 KG Querying (Graph)
In this strategy, a prompt is used to provide the LLM with the schema of the KG. The LLM is then instructed to generate a syntactically correct GQL statement that can answer the input textual query. Once generated, this GQL statement is executed and the results are further interpreted by the LLM to generate the answer.

### v2

@@ Abstract
Recent advances in Large Language Models (LLM) and Retrieval-Augmented Generation (RAG) techniques have improved data processing in network management. However, existing RAG methods like VectorRAG and GraphRAG struggle with the complexity and implicit nature of semi-structured technical data, leading to inefficiencies in time, cost, and retrieval. This paper introduces FastRAG, a novel RAG approach for semi-structured data. FastRAG proposes chunk sampling, schema learning, and script learning to extract and structure data without submitting entire data sources to the LLM. It integrates text search with knowledge graph (KG) querying to improve accuracy. The evaluation results demonstrate that FastRAG provides accurate question answering while improving up to $90\%$ in time and $85\%$ in cost compared to GraphRAG.
@@ I Introduction
Efficiently processing and understanding semi-structured data is crucial to improving network management tasks. As networks expand, the diversity of platforms and the sheer volume of data introduce significant challenges, including operational inefficiencies and delays in issue resolution. Traditional tools for processing network data [1, 2, 3] provide some utility, yet they fail to extract and leverage the full range of information embedded in semi-structured data formats such as logs and configurations. These limitations often hinder the ability to correlate, analyze, and act on insights, particularly in complex environments that involve heterogeneous devices and platforms.
Disaggregated implementations of network services further exacerbate the task of correlating semi-structured data from different vendors. This fragmentation not only complicates the parsing and normalization of data but also makes it challenging to derive actionable insights across varying formats. Although natural language processing (NLP) and machine learning (ML) algorithms have been developed to address these challenges [4, 5], they often lack the flexibility needed to handle the diversity of data types and formats encountered in real-world network environments.
@@
The emergence of Large Language Models (LLMs) has introduced new possibilities for processing and understanding natural language, showcasing their potential in improving network management tools by automating complex analytical tasks. However, using pre-trained models to generate domain-specific and coherent responses remains challenging. Techniques like Retrieval-Augmented Generation (RAG) [6] have been introduced to enhance LLMs by integrating retrieval methods, with VectorRAG providing context retrieval from textual documents based on semantic similarity [7]. Although effective in many scenarios, these traditional RAG systems have limitations, such as loss of critical contextual information due to uniform fragmentation of documents [8].
GraphRAG [9], a more advanced approach, addresses some of these limitations by leveraging knowledge graphs (KG) to organize information extracted from source documents. Despite its advantages, GraphRAG struggles with queries for accurate information or queries lacking explicit entities. HybridRAG [8] further refines these methods by combining the strengths of VectorRAG and GraphRAG, improving the precision in retrieving relevant information for LLM.
@@
Despite these advancements, existing RAG techniques do not meet the unique challenges posed by network data. The reliance on embedding vectors for context retrieval often underperforms with semi-structured technical data, where implicit information is frequently encoded in domain-specific keywords. Furthermore, current RAG systems rely heavily on LLMs to extract structured information from source documents by processing them chunk by chunk, leading to increased computational costs and inefficiencies, critical drawbacks when dealing with the large-scale, high-frequency data generated in modern networks.
To address these issues, we propose FastRAG, a novel RAG approach designed to process large volumes of semi-structured data in a cost-effective way. FastRAG addresses the specific needs of network management by offering a system optimized to extract actionable insights from diverse data sources while minimizing latency and cost. The system introduces innovative methods, including schema learning and script learning for structured data processing, an algorithm to select representative chunks for learning prompts, and a hybrid retrieval approach that integrates text search with graph querying using a query language. These methods enable FastRAG to deliver improved accuracy and efficiency, making it suitable for network management tasks such as troubleshooting, configuration auditing, and network analysis.
@@
The remainder of this paper is organized as follows. Section II reviews related work and discusses the motivation behind the development of FastRAG. Section III presents the detailed design of the FastRAG system, including its key components and techniques. The evaluation of FastRAG is discussed in Section IV. Finally, Section V outlines the system’s limitations and concludes the paper.
@@ II Related Work and Motivation
The integration of LLMs into network management has opened new possibilities, particularly in automating network configuration tasks. Recent studies have explored the potential of LLMs to generate network configurations [10, 11, 12], detect anomalies [13], and convert textual descriptions into formal specifications [14, 15]. These contributions represent the most frequent applications of LLMs in network management, aiming to reduce manual labor and error risks. The work proposed in this paper is complementary to these approaches, as they may rely on RAG systems to retrieve network-specific information.
@@
The application of RAG within network management has been less explored. RAGLog [16] introduces a system for detecting anomalies using a combination of vector database and LLM, supporting raw log data and various log sources without extensive pre-processing. Although promising for its adaptability, RAGLog faces resource consumption and latency challenges, which could limit its scalability in larger network data. Telco-RAG [17] targets the processing of 3GPP documents. A query enhancement pipeline demonstrates improved accuracy when handling complex questions related to the telecommunication domain.
@@
More generally, VectorRAG, the predominant approach in RAG systems, computes vector embeddings for fixed-size chunks and uses semantic similarity to retrieve information relevant to the input query. However, this method struggles with the nature of network data, where domain-specific keywords have different meanings in their context. GraphRAG [9] is a more advanced approach that integrates KGs with text chunking and embedding. The LLM extracts entities and relationships and summarizes communities detected in the KG, providing a more structured retrieval (also using semantic similarity). Although GraphRAG is efficient in summarizing and reporting tasks, it falls short in delivering valuable answers when the query includes specific values, such as names or types, that need to be matched precisely in the retrieval. For example, to the typical network management question ”What is the IP address of interface Ethernet0 on router r1?” GraphRAG would either not be able to answer the question, or provide a summary of several pieces from configuration files that refer to the mentioned interface. HybridRAG [8] combines GraphRAG and VectorRAG. Although this hybrid approach improves the accuracy and contextual relevance of responses, it inherits the limitations of its base approaches. Another approach uses entity-relationship extraction to create a KG capturing the important information of the source data. The KG is then searched as a database by translating the textual query into a database query. Although this allows for accurate retrieval, it can be impractical given the imperfect parsing obtained with LLMs [18].
@@
In contrast, FastRAG extracts entities and their properties from the source documents while mapping each entity to specific lines of the original text. This method enables accurate retrieval using a KG while allowing text search to handle vague queries and compensate for imperfect entity extraction. Unlike semantic similarity, the text search matches text based on the exact wording or structure. KG and text searches are performed using queries on a single KG implemented in a graph database. Unlike other methods, FastRAG avoids costly processing of all source data through LLMs by generating JSON schemas and Python code for data structuring and parsing. While demonstrated on logs and configuration data, this approach can be applied to other semi-structured data, such as playbooks and alarms. To our knowledge, this is the first RAG system to rely solely on code generation for data processing.
@@ III FastRAG Design
We designed a hybrid approach for both information extraction and retrieval. Instead of using indiscriminate chunking or a user-defined schema, our method converts the data into a simple, automatically generated JSON structure while maintaining a link to the original data in a KG [19]. Our RAG also relies on LLMs to interact with the KG, utilizing query generation to retrieve relevant information.
Given the extensive and continuously updating nature of network data, minimizing processing time and cost is essential. To achieve this, we leverage prompt engineering findings that indicate that LLMs are more effective at generating code rather than directly extracting information into structured formats. Therefore, we introduce schema learning and script learning, where we sample the source data and iteratively prompt the LLM to: (i) generate JSON schemas to structure the information, and (ii) generate Python scripts to parse the source data and extract entities.
Figure 1 shows the FastRAG data processing pipeline, illustrating chunk sampling, schema learning, and script learning. Each of these methods is described below, followed by a description of the information retrieval process.
Figure 1: FastRAG data processing architecture
@@ III-A Chunk sampling
To address LLM input size constraints, data is usually partitioned into uniformly sized, overlapping chunks and submitted via batched or sequential queries. However, this approach incurs significant latency and computational cost. To address this, we exploit the repetitive formatting present in semi-structured data, selecting a representative set of chunks, or sample chunks, that effectively covers the syntax of the source data [13]. The chunk sampling process involves two key procedures: keyword extraction and chunk selection.
@@ III-A1 Keyword extraction
We draw inspiration from NLP techniques to identify the most meaningful terms in the source data [4]. The process begins with the preprocessing of the text, where the punctuation is removed, and the content is tokenized into individual words. Given that English is the predominant language used in most semi-structured network data, it is adopted for this process. After preprocessing, the text is split into lines, and a matrix representation is created based on word frequency, where each row corresponds to a line, and each column to a unique term from the entire text. The matrix, which reflects the frequency of each term in each line, serves as the input for the K-means clustering algorithm. This algorithm groups the lines into $n_{c}$ groups based on the similarity of their frequency patterns. The $n_{t}$ terms closest to the centroids are selected within each group as keywords. The output is a set of keywords that encapsulate the primary terms present in the text corpus.
@@ III-A2 Samples selection
We developed an algorithm that selects the smallest set of chunks that contains all keywords. However, merely adding chunks until all keywords are covered does not guarantee the selection of the (near) minimum number of sample chunks. To address this, the algorithm selects chunks based on their entropy, providing an estimate of the diversity introduced by each chunk.
The procedure (see Algorithm 1) begins with preprocessing of the chunks (as in the previous step) where only the extracted keywords are retained. The preprocessed chunks are then used to compute Term Frequency-Inverse Document Frequency (TF-IDF) vectors, representing the importance of terms within each chunk. The Shannon entropy for each chunk is then calculated to gauge the informational diversity to which it contributes. The algorithm iteratively selects a subset of chunks to maximize term coverage, focusing on those that introduce new terms and the most diversity, until the desired coverage threshold - fixed at 1 in all cases - is reached.
@@
The number of sample chunks selected is directly influenced by the number of keywords extracted, which can be adjusted by varying the parameters $n_{c}$ (number of clusters) and $n_{t}$ (number of terms). This process is discussed in more detail in Section IV.
@@ III-B Schema learning
Building on the selected sample chunks, we develop a prompt strategy that guides the LLM in identifying entity types and their properties, focusing on schema extraction rather than specific entities. The process (Figure 3) begins with the first sample chunk, submitted to the LLM to identify and structure the types of entities and their properties in a JSON schema. The JSON schema is refined in subsequent prompts by submitting new chunks to the LLM. To avoid complexity, the LLM is instructed to limit high levels of nesting and refrain from creating objects with an excessive number of properties.
After each prompt, the JSON schema is validated. If the schema contains errors, the LLM is asked to correct them based on the error message provided. Each prompt call can be retried up to $N=4$ times.
Two types of objects are extracted from the finalized schema, as illustrated in Figure 2. The first is the Step 1 schema, which is a JSON schema that includes only the level-1 entity types -referred to as sections. Each entity type includes a description [^1] Not shown in the Figure and a string property containing the source data lines that define the objects that belong to that section. The second, Step 2 schema, is a map where each section is associated with an array where objects correspond to the entire schema of that section. This division into Step 1 and Step 2 schemas allows for a more structured approach to data extraction, as discussed below.
@@ III-C Script learning
Similar to the approach used in schema learning, the script learning prompt strategy (Figure 3) begins by instructing the LLM to generate parsing functions based on the first sample chunk. The LLM is then prompted iteratively to refine the previous function code by submitting new sample chunks. To ensure both the syntactic and functional correctness of the generated code, the prompt includes a verification step, where the code is executed with the sample chunk used to generate it. If the code contains errors, the LLM is asked to correct them based on the feedback message. Similarly, each prompt call can be retired up to N times.
In Step 1, the prompt instructs the LLM to map each line in the source data to its corresponding section according to the schema. Once the function is generated, it is used to parse the entire source data and divide it into sections. After that, each section is further split into fixed-size chunks, and the sampling process is repeated for each section. At this stage, each section has its schema (contained in the schema object in Step 2) and a set of sample chunks, enabling it to be processed independently, following the same method as in Step 1.
In Step 2, script learning is utilized to generate a specific parsing function for each section. Once these functions are generated, the data within each section is processed using the corresponding parsing function. Processing sections independently enables the LLM to concentrate more effectively on each entity type and also allows for a more targeted refinement of the sample chunks by re-executing the sampling process for each section.
@@ III-D Information Retrieval
The extraction process results in a JSON object where each entity is represented, and within this object, an input_data property contains the lines from the source data that define the entity. This JSON structure serves as the foundation for the creation of KGs and the retrieval of information.
@@ III-D1 KG Creation
Each entity identified in the JSON object is inserted as a node within the KG, the entity type serving as the node’s label. Simple-type properties of the entity are directly assigned as properties of the node, while properties that are themselves objects are inserted as child nodes, connected to their parent node. Furthermore, for each line within the input_data property of each entity, a corresponding node is created and linked to the parent entity node. These input data nodes are used for text searches that utilize NLP methods [20].
@@ III-D2 KG Querying (Graph)
A prompt is used to provide the LLM with the KG schema. The LLM is then instructed to generate a syntactically correct query statement that can answer the user input. The generated query is executed, and the results are interpreted by the LLM to generate the final answer.